write y=7/2x-6 in standard form.show your work justify each steo
Answer by Alan3354(69443)

y=7/2x-6
---------
I'll assume you mean (7/2)x and not 7/(2x):
Multiply by 2
2y = 7x - 12
2y + 12 = 7x
7x - 2y = 12